OpenStudy (anonymous):

|6y-2|+4<22 Please solve, thanks

OpenStudy (anonymous):

|6y-2| < 22 - 4 |6y - 2| < 18 6y - 2 = +/- 18 6y = 20 y = 10/3 OR 6y = -16 y = -8/3 so y = 10/3, -8/3

OpenStudy (anonymous):

fucin smart